# Feasibility of Non-Line-of-Sight Integrated Sensing and Communication at mmWave

### v1

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ II ISAC Setup
$d_{\text{res}}$ $=\frac{c_{0}}{2\Delta fN}=0.79~{}\text{m}\;.$ (4)
@@ III-A Impact of Time Domain Holes
In standard sensing operations, we process the whole CSI matrix $\mathbf{H}$ corresponding to each frame to maximize the processing gain and to achieve the speed resolution as in Eq. (3). However, since the gNB transmits in a TDD pattern, UL symbols do not contain useful information for radar processing and are discarded (set to zero) before computing the periodogram. Using TDD, each UL/DL pattern is repeated $T_{f}/T_{\text{TDD}}=R$ times per frame. With the parameters from Section II, this amounts to $R=8$ TDD patterns, as shown in Fig. 1. Setting the UL parts to zero acts as an additional windowing effect on $\mathbf{H}$ and introduces spectral replicas in the speed domain, as shown in Fig. 2. This “on/off” windowing is described by a discrete rectangle function (Dirichlet window) convolved with a train of Dirac deltas in the time domain
$w_{\text{TDD}}(k)$ $=\text{rect}\left(\frac{k+M_{\text{UL}}/2}{M_{\text{DL}}\cdot T_{s}}\right)\ast\sum_{i=0}^{7}\delta\left(k-i\cdot T_{\text{TDD}}\right)\;.$ (7)
Its Fourier transform is a cardinal sine (Dirichlet Kernel) function multiplied by a train of Dirac deltas, which is convolved with the target in the speed (Doppler) domain. Therefore, after windowing and computing the periodogram, each target experiences replicas spaced by $R\cdot v_{\text{res}}=4.4$ m/s in speed due to the frequency of the windowing in time.
@@ III-B CSI Processing Techniques
Processing only the first DL sequence with $M_{\text{DL}}=104$ symbols would avoid replicas. However, per Eq. (3), this reduces the speed resolution to $5.88$ m/s, which is unacceptable for most use cases.
Alternatively, the CSI matrix can be downsampled in time by selecting only every $J$-th OFDM symbol to avoid UL parts. This approach is effective in removing the replicas but reduces the unambiguous speed due to the decimation factor $J$. Further, the processing gain is decreased due to the reduced number of processed symbols, leading to a lower SNR.
@@ V Measurement Results
The detection rate was estimated for the 10 strongest NLOS returns, comparing them with the LOS components. With the focus on intrusion detection, it is sufficient to detect the presence of the intruder once and raise an alarm accordingly. A possible approach for discriminating the presence of actual targets from false alarms is to verify that NLOS peaks are continuously detected within a window of consecutive updates. In our measurements, an overall detection rate of up to 67% was obtained with $K=8$ frames and stride $V=2$ frames.
@@ VI Conclusion
Figure 6: Moving average of NLOS detections over a 0.5 s time window, using exponential threshold and clutter removal. Periodograms were generated using CSI matrices from $K=8$ frames.

### v2

@@ Abstract
© 20XX IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ II ISAC Setup
$d_{\text{res}}$ $=\frac{c_{0}}{2N\Delta f}=0.79~{}\text{m}\;.$ (4)
@@ III-A Impact of Time Domain Holes
In standard sensing operations, we process the whole CSI matrix $\mathbf{H}$ for each frame to maximize the processing gain and to achieve the speed resolution as in Eq. (3). However, since the gNB transmits in a TDD pattern, UL symbols do not contain useful information for radar processing and are discarded (set to zero) before computing the periodogram. Using TDD, each UL/DL pattern is repeated $T_{f}/T_{\text{TDD}}=R$ times per frame. With the parameters from Section II, this amounts to $R=8$ TDD patterns, as shown in Fig. 1. Setting the UL parts to zero acts as an additional windowing effect on $\mathbf{H}$ and introduces spectral replicas in the speed domain, as shown in Fig. 2. This “on/off” windowing is described by a discrete rectangle function, with a width of $M_{\text{DL}}$ symbols, convolved with a train of Dirac deltas in the time domain
$w(t)$ $=\negthickspace\negthickspace\sum_{k=0}^{M_{\text{DL}}-1}\negthickspace\negthickspace\delta(t-kT_{s})\ast\negthickspace\negthickspace\sum_{k=-\infty}^{\infty}\negthickspace\delta\left(t-kT_{\text{TDD}}+\frac{M_{\text{DL}}T_{s}}{2}\right)\!.$ (7)
The resulting point spread function is the Fourier transform of (7), consisting of a Dirichlet Kernel multiplied by a train of Dirac deltas, that is convolved with the targets’ contributions in the speed domain. As the TDD pattern repeats $R$ times within the time aperture $MT_{s}$, the point spread function shows contributions in the speed domain spaced by
[…]
This is the source of the speed replicas that can be again observed to appear at $4.4$ m/s distance in Fig. 2.
@@ III-B CSI Processing Techniques
Processing only the first DL sequence with $M_{\text{DL}}=104$ symbols would avoid replicas. However, per Eq. (3), this reduces the speed resolution to $5.88$ m/s, which is unacceptable for most use cases. Alternatively, the CSI matrix can be downsampled in time by selecting every $J$-th OFDM symbol such that UL parts are avoided. This approach is effective in removing the replicas but reduces the unambiguous speed by a factor of $J$, as well as the processing gain due to the reduced number of processed symbols, leading to a lower SNR.
@@ V Measurement Results
Detection rate was estimated for the 10 strongest NLOS returns, comparing them with the LOS components. With the focus on intrusion detection, it is sufficient to detect the presence of the intruder once and raise an alarm accordingly. A possible approach for discriminating the presence of actual targets from false alarms is to verify that NLOS peaks are continuously detected within a window of consecutive updates. In our measurements, an overall detection rate of up to 67% was obtained with $K=8$ frames and stride $V=2$ frames.
@@ VI Conclusion
Figure 6: Moving average of NLOS detection rate over a 0.5 s time window, using exponential threshold and clutter removal. Periodograms were generated using CSI matrices from $K=8$ frames.